Exercise 2-9 (Part Level Submission) Usonic, Inc., has collected the following information on its cost of electricity: January February March April May June July August September October November December Machine Hours 629 700 494 427 448 304 373 558 573 274 428 200 Total Electricity Costs $280 $320 $272 $208 $237 $182 $180 $246 $260 $174 $220 $120 Your answer is incorrect. Try again. Compute the total fixed cost of electricity. (Round answer to 2 decimal places, e.g. 15.25.) Fixed cost
Answer #1
Using High low method , we can split a mixed cost into its fixed and variable components.
The cost formula used in this method is as under,
y = a + bx
y = Total Cost
a = Fixed cost
b = Variable cost per unit
x = number of units
The formula to calculate b i.e.variable cost per unit is as under,
Variable cost per unit = [y2 - y1]/[x2 - x1]
y2 = total cost at highest level of activity = $320
y1 = total cost at lowest level of activity = $120
x2 = the number of machine hours at highest level of activity = 700
x1 = the number of machine hours at lowest level of activity = 200
Variable cost per unit = [$320 - $120]/[700 - 200] = $0.40 per machine hour
Fixed cost i.e.a = y - bx
Fixed cost i.e.a = $320 - ($0.40 *700) = $40